How do you evaluate expressions involving radicals? 4,square root symbol over -20736
Answer by stanbon(75887)

How do you evaluate expressions involving radicals? 4,square root symbol over -20736
-----------------
The 4th root of 20736 is 12
-------------------
How do you get that without a calculator?
You express 20736 as a product of primes:
20736 = 2*10368
=2*2*5184
=2*2*2*2592
=2*2*2*2*1296
=2*2*2*2*2*648
=2^5*2*324
=2^5*2*2*162
=2^7*2*81
=2^8*3^4
-------------
4th root or 2^8*3^4 = 2^2*3 = 12
====================
Cheers,
Stan H.